OpenStudy (anonymous):

15. Simplify the expression below. Show your work. –(–7y + 12)

OpenStudy (anonymous):

You can treat the negative in front like a negative 1: $$-1 \cdot (-7y + 12)$$ Simply distribute the negative into the parenthesis: $$(-7y \cdot -1) + (12 \cdot -1)$$ $$(+7y) + (-12)$$ $$7y -12$$

OpenStudy (anonymous):

In another sense, the negative just makes all of the signs (+ or -) switch within the parenthesis.